OpenStudy (anonymous):

Write an equation to represent the problem. The area of a triangle is equal to one half times the base times the height. Use A for area, b for base and h for height. A. A = 1/2 b + h B. A = 1/2 (b + h) C. A = (1/2) bh

OpenStudy (lilg132):

c

OpenStudy (anonymous):

can you tell me how you got it please?

OpenStudy (anonymous):

because multiplication is commutative It doesn't matter what order thy are multiplied in

OpenStudy (lilg132):

A = (1/2) bh when expanded looks like this: 1/2b x h = A or 1/2 x b x h = A

OpenStudy (anonymous):

ok

OpenStudy (lilg132):

no medal?

OpenStudy (lilg132):

no medal given